# Transaction Fee Market Design for Parallel Execution

### v1

@@ Abstract.
Given the low throughput of blockchains like Bitcoin and Ethereum, scalability — the ability to process an increasing number of transactions — has become a central focus of blockchain research. One promising approach is the parallelization of transaction execution across multiple threads. However, achieving efficient parallelization requires a redesign of the incentive structure within the fee market. Currently, the fee market does not differentiate between transactions that access multiple high-demand resources versus a single low-demand one, as long as they require the same computational effort. Addressing this discrepancy is crucial for enabling more effective parallel execution.
@@
In this work, we aim to bridge the gap between the current fee market and the need for parallel execution by exploring alternative fee market designs. To this end, we propose a framework consisting of two key components: a Gas Computation Mechanism (GCM), which quantifies the load a transaction places on the network in terms of parallelization and computation, measured in units of gas, and a Transaction Fee Mechanism (TFM), which assigns a price to each unit of gas. We also introduce a set of desirable properties for a GCM, present multiple candidate mechanisms, and evaluate them against the properties. One promising candidate emerges: the weighted area GCM. Notably, this mechanism can be seamlessly composed with existing TFMs, such as EIP-1559. While our exploration primarily focuses on the execution component of the fee, which directly relates to parallel execution, we also outline how it could be integrated with fees associated with other factors, such as storage and data bandwidth, by drawing a parallel to a multi-dimensional fee market.
1. Introduction
@@
Scalability — the ability to process more transactions efficiently — has become a central focus in blockchain research, especially given the low throughput of many existing networks. Ethereum, for example, is constrained by its single-threaded execution model, limiting transaction throughput. One promising way to enhance scalability is by parallelizing transaction execution across multiple threads, taking advantage of the multi-core processors common in modern hardware. However, achieving the full efficiency gains of parallel execution requires rethinking fee market design to better account for resource contention and scheduling constraints.
@@
A transaction fee mechanism (TFM) is a core component of any blockchain protocol, determining which pending transactions should be processed and what users must pay for the privilege of having their transactions executed. Traditional fee mechanisms, like Bitcoin’s first-price auction, involve users submitting bids with their transactions, and the transactions with the highest bids per computation are included in the next block. Ethereum initially used a similar purely first-price auction-based model but switched to the more sophisticated EIP-1559 mechanism in 2021, which introduces a fluctuating base fee based on network demand, aiming to improve incentive compatibility and reduce price volatility (Buterin et al., 2024). Most of the existing literature on transaction fee mechanisms focuses on settings where transactions are executed sequentially and therefore does not account for resource contention, which is crucial in parallel execution environments.
@@
Thus, viewed in isolation, these traditional TFMs are ill-suited to the complexities introduced by parallel transaction execution. They price transactions solely based on their computational cost, without distinguishing between those that access a single resource and those that interact with multiple, potentially contested resources. This pricing model works in a single-threaded environment but fails to capture the nuances of parallel execution, where transactions may impose vastly different constraints on resource scheduling. Transactions that touch multiple high-contention resources can introduce bottlenecks, while those interacting with isolated resources are far easier to schedule efficiently.
@@
Ethereum has yet to adopt parallel execution, but several blockchains, such as Solana, Aptos, and Sui, already employ parallel transaction processing (Aptos Labs, 2023; Monad Labs, 2023; Sei Labs, 2023; Solana Labs, 2023; Mysten Labs, 2023). However, many of these networks have yet to implement fee models that fully account for the challenges of parallel execution. Sui and Solana have introduced fee markets tailored to parallelization, but these mechanisms require users to engage in first-price auctions for congested resources (Lostin, 2025; Mueller, 2025). As a result, these fee markets demand a high level of sophistication from users to effectively optimize their fee settings and are also not incentive-compatible. The requirements for an effective fee market that is suitable for parallel execution and the design of such a market have so far remained unresolved.
@@
  • Our analysis identifies a promising candidate: the weighted area GCM. Importantly, this mechanism is not only an effective GCM, achieving a large subset of the outlined properties, but it is equally important that it can be easily integrated with existing TFMs, such as EIP-1559, inheriting their properties.
2. Model
We consider a universe consisting of several stateful resources (e.g., user accounts, storage addresses of smart contracts). Each resource can be thought of as a system global variable. We write $\mathcal{R}$ for the set of resources. For analysis purposes, we assume that $\mathcal{R}$ is infinite. A transaction is a sequence of elementary instructions performing computation and interacting with the resources. Some of these operations access a given target resource (e.g., read its value, write to it). For simplicity, we assume that the following are known in advance and supplied with the transaction:
  • The non-empty set of resources $R\subseteq\mathcal{R}$ the transaction accesses, or an overestimate of it. [^1] In Ethereum, transaction accesses are generally not known in advance. Instead, transactions can execute arbitrary logic (constrained by a maximum amount of computational effort) and their execution depends on the blockchain’s state at the time of execution. Ethereum currently supports optional access lists that allow transactions to specify their accesses (Buterin and Swende, 2020). This optional list could be made mandatory to provide an overestimate of accesses. Additionally, it is worth noting that in other blockchains (e.g., Solana (Labs, 2025b) and Sui (Labs, 2025a)), an overestimate of accesses is typically known in advance.
@@ 2. Model
Transactions execute concurrently, but atomically, meaning that the overall effect of executing a batch of transactions should also be achievable by a sequential, single-core execution. For simplicity, we restrict to concurrent schedules following a simple lock-based execution policy: each resource has a lock associated with it; whenever a thread wants to execute a transaction, it first locks all required resources, then executes the transaction, and then releases the locks. We assume that acquiring (and similarly releasing) the required locks happens simultaneously and takes no additional time. These simplifications have a desirable side-effect: $t$ and $R$ for each transaction now uniquely determine the set of admissible concurrent schedules, allowing us to ignore other details about the transactions:
@@
A transaction tx is specified through a tuple $(t,R)$, where $t>0$ denotes the time required to execute the transaction and $R\neq\varnothing$ represents the set of resources the transaction demands, which are locked for the duration of the transaction. We will sometimes write $t(\textit{tx})$ and $R(\textit{tx})$ for $t$ and $R$ respectively.
@@
In Figure 1, we illustrate a sample transaction $\textit{tx}\simeq(3,\{r_{2},r_{3}\})$. Transaction tx thus takes 3 units of time to execute and utilizes resources $r_{2}$ and $r_{3}$. We illustrate this by a rectangle of corresponding length (i.e., time) and width (i.e., resources). Throughout, we will illustrate transactions in this manner to aid in visualizing concepts and results. Note that we will always represent transactions as rectangles (i.e., they use consecutive resources). In reality, this is, of course, not the case. Importantly, all our results also hold in the more general setting where a transaction can use any subset of resources.
time3r1r2r3r4resources
Figure 1. Illustration of a sample transaction $\textit{tx}\simeq(3,\{r_{2},r_{3}\})$.
@@
A blockchain is a sequence of blocks comprising bundles of transactions: $B_{1},\dots,B_{k}$. The system starts in some predetermined initial state. The blocks are executed in order starting from the oldest (the genesis block $B_{1}$), successively changing the system’s state. For simplicity, we assume no cross-block parallelism, so the execution of a block only starts after the previous block’s execution has been completed. However, the execution of transactions inside a block happens concurrently. The system’s state after executing $B_{k}$ is the current system state. Users wanting to change the system’s state compete for inclusion in the next block $B_{k+1}$ and have to pay a fee if successfully included. Desirably, the fee should be higher for more complex transactions and higher during periods of high demand due to limited block space. These requirements are typically decoupled and ensured through different means:
  1. (1)
@@
    A transaction’s complexity is quantified in units of gas: [^3] For Ethereum, the unit is called wei. the more complex a transaction is, the more gas it consumes. Gas encompasses multiple components such as execution, storage space, and data bandwidth. For our purposes, we will only be concerned with execution gas. [^4] From this point forward, gas will refer specifically to execution gas. Currently, the execution gas acts as a proxy for the execution time $t$, [^5] Throughout we will assume that execution gas exactly corresponds to time. but this need not be the case, as we will demonstrate in our paper.
  2. (2)
@@
    The fair competition for block space is ensured through a transaction fee mechanism (TFM): users submit transactions they would like to be included in the next block together with bids of what they would be willing to pay per unit of gas. Importantly, block space is limited, i.e., there is limited space for transactions. The mechanism then determines the set of transactions to be included in the block, together with a price per unit of gas to be paid by each included transaction (potentially not the same for all transactions). [^6] There are several other components of a TFM, but for the level of detail we need here, this suffices.
@@
A gas computation mechanism (GCM) takes as input a set of transactions $T$ to be included in a block and determines in a deterministic manner the amount of gas consumed by each transaction $\textit{tx}\in T$, written $\textsf{gas}_{T}(\textit{tx})$.
Currently deployed GCMs associate a fixed, predetermined gas consumption with each transaction, independent of the specific resources accessed by the transaction, making them unsuitable for a parallel execution environment. In particular, this is true for Ethereum’s current GCM:
@@
Our goal is to ensure that fees accurately reflect the parallelizability of transactions. Therefore, the gas consumption calculated for a transaction should depend on the set of resources it accesses and may also be influenced by factors external to the transaction itself (like interactions between transactions).
3. GCM Properties
@@
We begin with two natural monotonicity properties, one for resources and one for time. First, a transaction that requires a subset of the resources used by another transaction, while taking the same amount of time, should consume no more gas (Properties 1 and 5). Similarly, a transaction that requires no more execution time than another, assuming both involve the same set of resources, should consume no more gas (Properties 2 and 5).
timer1r2r3r4resources
timer1r2r3r4resources
Figure 2. Illustration of Property 1, where $T=\emptyset$,    represents a sample $\textit{tx}_{1}$ and    represents a sample $\textit{tx}_{2}$.
timer1r2r3r4resources
timer1r2r3r4resources
Figure 3. Illustration of Property 2, where $T=\emptyset$,    represents a sample $\textit{tx}_{1}$ and    represents a sample $\textit{tx}_{2}$.
timer1r2r3r4resources
timer1r2r3r4resources
Figure 4. Illustration of Property 4, where $T=\emptyset$, $T_{1}=\{\textit{tx}_{1},tx_{2}\}$, and $T_{2}=T_{1}\cup\{\textit{tx}_{3}\}$. Here,    represents a sample $\textit{tx}_{1}$,    represents a sample $\textit{tx}_{2}$, and    represents a sample $\textit{tx}_{3}$.
timer1r2r3r4resources
timer1r2r3r4resources
Figure 5. Illustration of Property 5, where $T=\emptyset$,    represents a sample $\textit{tx}_{3}$, while    represents a sample $\textit{tx}_{1}$ and    represents a sample $\textit{tx}_{2}$. The darker shaded areas indicate when a transaction operates on a resource.
Property 1 (Resource Sensitivity).
Given a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},R_{1})$ and $\textit{tx}_{2}\simeq(t_{2},R_{2})$, both not in $T$, such that $t_{1}=t_{2}$ and $R_{1}\subseteq R_{2}$:
@@
Property 2 (Time Sensitivity).
Given a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},R_{1})$ and $\textit{tx}_{2}\simeq(t_{2},R_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $R_{1}=R_{2}$:
@@
Property 3 (Resource-Time Sensitivity).
Given a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},R_{1})$ and $\textit{tx}_{2}\simeq(t_{2},R_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $R_{1}\subseteq R_{2}$:
@@
Intuitively, if $\textit{tx}_{1}\lesssim\textit{tx}_{2}$, by which we mean $t(\textit{tx}_{1})\leq t(\textit{tx}_{2})$ and $R(\textit{tx}_{1})\subseteq R(\textit{tx}_{2})$, then $\textit{tx}_{1}$ should cost no less than $\textit{tx}_{2}$. However, an attentive reader will observe that the former two are collectively equivalent to the latter (the proof and all subsequent omitted proofs can be found in the appendix):
@@
Property 3 holds if and only if Properties 1 and 2 hold.
Let us now take a moment to briefly evaluate why Properties 1, 2 and 3 are not merely intuitive, but their violation can lead to harmful consequences and misaligned incentives: suppose $\textit{tx}_{1}\lesssim\textit{tx}_{2}$ and $T$ is a set of transactions containing neither of the two. If $\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}(\textit{tx}_{1})>\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}(\textit{tx}_{2})$, a user intending to submit $\textit{tx}_{1}$ might instead pad $\textit{tx}_{1}$ with unnecessary instructions and declare a larger access list to decrease the gas consumption. This would paradoxically reduce the gas usage and, assuming a reasonable TFM is used to compute transaction fees, also lower the transaction fee.
For any of the three properties above, we say that the property is strictly satisfied if for $\textit{tx}_{1}\not\simeq\textit{tx}_{2}$ the conclusion inequality holds strictly. Note that properties holding strictly are even more desirable with respect to the reasoning above: replacing a transaction with a “larger” one is then not only no better but actively worse. Unsurprisingly, Lemma 3.1 also holds for the strict versions of the properties:
@@
Property 3 holds strictly if and only if Properties 1 and 2 hold strictly.
Next, we introduce another desirable monotonicity property, this time with a different emphasis: if two sets of transactions satisfy $T_{1}\subseteq T_{2}$, the transactions in $T_{1}$ should collectively consume no more gas than those in $T_{2}$ (Properties 4 and 5). To formalize this, given a set of transactions $T$ and a subset $T^{\prime}\subseteq T$, write $\textsf{gas}_{T}(T^{\prime}):=\sum_{\textit{tx}^{\prime}\in T^{\prime}}\textsf{gas}_{T}(\textit{tx}^{\prime})$ for the total gas consumed by the transactions in $T^{\prime}$ when included in the set $T$ that constitutes a block.
@@
Similarly to before, we say that Property 4 holds strictly if for $T_{1}\neq T_{2}$ the inequality in the conclusion holds strictly, which is desirable for reasons similar to those discussed above.
We now move on to more complex properties. Given two transactions $\textit{tx}_{1}\simeq(t_{1},R_{1})$ and $\textit{tx}_{2}\simeq(t_{2},R_{2}),$ a transaction $\textit{tx}_{3}$ is their sequential composition (or, more simply, concatenation), if it executes the steps of $\textit{tx}_{1}$ followed by the steps of $\textit{tx}_{2}$. Note that, in this case $\textit{tx}_{3}\simeq(t_{1}+t_{2},R_{1}\cup R_{2})$. Our next property states that the concatenation of two transactions should consume no less gas than submitting them individually (Properties 5 and 5). The two individual transactions perform the same actions as their concatenation, but not atomically, with no guarantee of their relative ordering or control over what happens between them. Naturally, enforcing atomicity and ordering limits the set of admissible concurrent schedules and requires resources to remain locked over longer continuous timespans. In particular, $\textit{tx}_{3}$ requires the resources in $R_{1}\cup R_{2}$ to be locked over a continuous span of $t_{1}+t_{2}$ units, while $\textit{tx}_{1}$ and $\textit{tx}_{2}$ submitted individually only require exclusive access to resources in $R_{1}$ for $t_{1}$ units and to resources in $R_{2}$ for $t_{2}$ units. Hence, the “bigger” transaction is at least as hard to schedule as its two constituent “parts” and should hence consume no less gas.
@@
We say that Property 5 holds strictly if the inequality in the conclusion holds strictly, which is again more desirable than the basic version of the property.

### v2

@@ Abstract
Given the low throughput of blockchains like Bitcoin and Ethereum, scalability — the ability to process an increasing number of transactions — has become a central focus of blockchain research. One promising approach is the parallelization of transaction execution across multiple threads. However, achieving efficient parallelization requires a redesign of the incentive structure within the fee market. Currently, the fee market does not differentiate between transactions that access multiple high-demand storage keys (i.e., unique identifiers for individual data entries) versus a single low-demand one, as long as they require the same computational effort. Addressing this discrepancy is crucial for enabling more effective parallel execution.
@@
In this work, we aim to bridge the gap between the current fee market and the need for parallel execution by exploring alternative fee market designs. To this end, we propose a framework consisting of two key components: a Gas Computation Mechanism (GCM), which quantifies the load a transaction places on the network in terms of parallelization and computation, measured in units of gas, and a Transaction Fee Mechanism (TFM), which assigns a price to each unit of gas. We additionally introduce a set of desirable properties for a GCM, propose several candidate mechanisms, and evaluate them against these criteria. Our analysis highlights two strong candidates: the weighted area GCM, which integrates smoothly with existing TFMs such as EIP-1559 and satisfies a broad subset of the outlined properties, and the time-proportional makespan GCM, which assigns gas costs based on the context of the entire block’s schedule and, through this dependence on the overall execution outcome, captures the dynamics of parallel execution more accurately.
keywords:
blockchain, transaction fee mechanism, parallel execution
category:
1 Introduction
@@
Scalability, the ability to process more transactions efficiently, has become a central focus in blockchain research, especially given the low throughput of many existing networks. Ethereum, for example, is constrained by its single-threaded execution model, limiting transaction throughput. One promising way to enhance scalability is by parallelizing transaction execution across multiple threads, taking advantage of the multi-core processors common in modern hardware. However, achieving the full efficiency gains of parallel execution requires rethinking the fee market design to better account for storage key contention and scheduling constraints. A storage key is a unique identifier (like an address label) for a specific data item in storage.
@@
A transaction fee mechanism (TFM) is a core component of any blockchain protocol, determining which pending transactions should be processed and what users must pay for the privilege of having their transactions executed. Traditional fee mechanisms, like Bitcoin’s first-price auction, involve users submitting bids with their transactions, and the transactions with the highest bids per computation are included in the next block. Ethereum initially used a similar purely first-price auction-based model but switched to the more sophisticated EIP-1559 mechanism in 2021, which introduces a fluctuating base fee based on network demand, aiming to improve incentive compatibility and reduce price volatility [11]. Most of the existing literature on transaction fee mechanisms focuses on settings where transactions are executed sequentially and therefore does not account for storage key contention, which is crucial in parallel execution environments.
@@
Thus, viewed in isolation, these traditional TFMs are ill-suited to the complexities introduced by parallel transaction execution. They price transactions solely based on their computational cost, without distinguishing between those that access a single storage key and those that interact with multiple, potentially contested storage keys. This pricing model works in a single-threaded environment but fails to capture the nuances of parallel execution, where transactions may impose vastly different constraints on storage key scheduling. Transactions that touch multiple high-contention storage keys can introduce bottlenecks, while those interacting with isolated storage keys are far easier to schedule efficiently.
@@
Ethereum has yet to adopt parallel execution, but several blockchains, such as Solana, Aptos, and Sui, already employ parallel transaction processing [4, 42, 56, 59, 45]. However, many of these networks have yet to implement fee models that fully account for the challenges of parallel execution. Sui and Solana have introduced fee markets tailored to parallelization, but these mechanisms require users to engage in first-price auctions for congested storage keys [39, 43]. As a result, these fee markets demand a high level of sophistication from users to effectively optimize their fee settings and are also not incentive-compatible. The requirements for an effective fee market that is suitable for parallel execution and the design of such a market have so far remained unresolved.
@@
  • Our analysis identifies two promising candidates. The weighted area GCM achieves a large subset of the outlined properties and, importantly, can be seamlessly integrated with existing TFMs, inheriting their properties. Complementing it, the time-proportional makespan GCM is designed to price execution costs in proportion to the total computational load of a block, allowing more accurate resource pricing at the block level.
2 Model
We consider a universe consisting of several stateful storage keys (e.g., user accounts, storage addresses of smart contracts). Each storage key can be thought of as a system global variable. We write $\mathcal{K}$ for the set of storage keys. For analysis purposes, we assume that $\mathcal{K}$ is infinite. A transaction is a sequence of elementary instructions performing computation and interacting with the storage keys. Some of these operations access a given target storage key (e.g., read its value, write to it). For simplicity, we assume that the following are known in advance and supplied with the transaction:
  • The non-empty set of storage keys $K\subseteq\mathcal{K}$ the transaction accesses, or an overestimate. [^2] In Ethereum, transaction accesses are generally not known in advance. Instead, transactions can execute arbitrary logic (constrained by a maximum amount of computational effort) and their execution depends on the blockchain’s state at the time of execution. Ethereum currently supports optional access lists that allow transactions to specify their accesses [13]. This optional list could be made mandatory to provide an overestimate of accesses. Additionally, it is worth noting that in other blockchains (e.g., Solana [35] and Sui [34]), an overestimate of accesses is typically known in advance.
@@ 2 Model
Transactions execute concurrently, but atomically, meaning that the overall effect of executing a batch of transactions should also be achievable by a sequential, single-core execution. For simplicity, we restrict to concurrent schedules following a simple lock-based execution policy: each storage key has a lock associated with it; whenever a thread wants to execute a transaction, it first locks all required storage keys, then executes the transaction, and then releases the locks. We assume that acquiring (and similarly releasing) the required locks happens simultaneously and takes no additional time. These simplifications have a desirable side-effect: $t$ and $K$ for each transaction now uniquely determine the set of admissible concurrent schedules, allowing us to ignore other details about the transactions:
@@
A transaction tx is specified through a tuple $(t,K)$, where $t>0$ denotes the time required to execute the transaction and $K\neq\varnothing$ represents the set of storage keys the transaction demands, which are locked for the duration of the transaction. We will sometimes write $t(\textit{tx})$ and $K(\textit{tx})$ for $t$ and $K$ respectively.
@@
In Figure 1, we illustrate a sample transaction $\textit{tx}\simeq(3,\{k_{2},k_{3}\})$. Transaction tx thus takes 3 units of time to execute and utilizes storage keys $k_{2}$ and $k_{3}$. We illustrate this by a rectangle of corresponding length (i.e., time) and width (i.e., storage keys). Throughout, we will illustrate transactions in this manner to aid in visualizing concepts and results. Note that we will always represent transactions as rectangles (i.e., they use consecutive storage keys). In reality, this is, of course, not the case. Importantly, all our results also hold in the more general setting where a transaction can use any subset of storage keys.
[…]
Figure 1: Illustration of a sample transaction $\textit{tx}\simeq(3,\{k_{2},k_{3}\})$.
@@
A blockchain is a sequence of blocks comprising bundles of transactions: $B_{1},\dots,B_{m}$. The system starts in some predetermined initial state. The blocks are executed in order, starting from the oldest (the genesis block $B_{1}$), successively changing the system’s state. For simplicity, we assume no cross-block parallelism, so the execution of a block only starts after the previous block’s execution has been completed. However, the execution of transactions inside a block happens concurrently. The system’s state after executing $B_{m}$ is the current system state. Users wanting to change the system’s state compete for inclusion in the next block $B_{m+1}$ and have to pay a fee if successfully included. Desirably, the fee should be higher for more complex transactions and higher during periods of high demand due to limited block space. These requirements are typically decoupled and ensured through different means:
  1. 1.
@@
    A transaction’s complexity is quantified in units of gas: [^4] For Ethereum, the unit is called wei. the more complex a transaction is, the more gas it consumes. Gas encompasses multiple components such as execution, storage space, and data bandwidth. For our purposes, we will only be concerned with execution gas. [^5] From this point forward, gas will refer specifically to execution gas. Currently, the execution gas acts as a proxy for the execution time tt, [^6] Throughout we will assume this approximation to the exact. but this need not be the case, as we will demonstrate in our paper.
  2. 2.
@@
    The fair competition for block space is ensured through a transaction fee mechanism (TFM): users submit transactions they would like to be included in the next block together with bids of what they would be willing to pay per unit of gas. Importantly, block space is limited, i.e., there is limited space for transactions. The mechanism then determines the set [^7] In practice, blocks are ordered, and transactions often compete for earlier positions (particularly in MEV settings). Following a common simplification in the TFM literature, we model blocks as sets of transactions. Accounting for intra-block ordering would require extending our model and constraining it to schedulers that honor the specified ordering. Our framework appears reasonably extensible to such settings, and we leave a detailed treatment to future work. of transactions to be included in the block, together with a price per unit of gas to be paid by each included transaction (potentially not the same for all transactions). [^8] There are several other components of a TFM, but for the level of detail we need here, this suffices.
@@
A gas computation mechanism (GCM) takes as input a set of transactions $T$ to be included in a block and determines in a deterministic [^9] We focus solely on deterministic mechanisms. This is common in blockchain fee markets, given the inherent difficulty of accessing true randomness on-chain (though notable approaches exist, such as VRFs). While randomized mechanisms can statistically improve over their deterministic counterparts, their ex-post behavior is often hard to predict, leading to higher fee uncertainty, an outcome generally undesirable in blockchain settings. manner the amount of gas consumed by each transaction $\textit{tx}\in T$, written $\textsf{gas}_{T}(\textit{tx})$.
Currently deployed GCMs associate a fixed, predetermined gas consumption with each transaction, independent of the specific storage keys accessed by the transaction, making them unsuitable for a parallel execution environment. In particular, this is true for Ethereum’s current GCM:
@@
Our goal is to ensure that fees accurately reflect the parallelizability of transactions. Therefore, the gas consumption calculated for a transaction should depend on the set of storage keys it accesses and may also be influenced by factors external to the transaction itself (like interactions between transactions).
3 GCM Properties
@@
We begin with two natural monotonicity properties, one for storage keys and one for time. First, a transaction that requires a subset of the storage keys used by another transaction, while taking the same amount of time, should consume no more gas (Properties˜1 and 5). Similarly, a transaction that requires no more execution time than another, assuming both involve the same set of storage keys, should consume no more gas (Properties˜2 and 5).
[…]
Figure 2: Illustration of Property˜1, where $T=\emptyset$,    represents a sample $\textit{tx}_{1}$ and    represents a sample $\textit{tx}_{2}$.
[…]
Figure 3: Illustration of Property˜2, where $T=\emptyset$,    represents a sample $\textit{tx}_{1}$ and    represents a sample $\textit{tx}_{2}$.
[…]
Figure 4: Illustration of Property˜4, where $T=\emptyset$, $T_{1}=\{\textit{tx}_{1},tx_{2}\}$, and $T_{2}=T_{1}\cup\{\textit{tx}_{3}\}$. Here,    represents a sample $\textit{tx}_{1}$,    represents a sample $\textit{tx}_{2}$, and    represents a sample $\textit{tx}_{3}$.
[…]
Figure 5: Illustration of Property˜5, where $T=\emptyset$,    represents a sample $\textit{tx}_{3}$, while    represents a sample $\textit{tx}_{1}$ and    represents a sample $\textit{tx}_{2}$. The darker shaded areas indicate when a transaction operates on a storage key.
Property 1 (Storage Key Monotonicity).
Given a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$, both not in $T$, such that $t_{1}=t_{2}$ and $K_{1}\subseteq K_{2}$:
@@
Property 2 (Time Monotonicity).
Given a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $K_{1}=K_{2}$:
@@
Property 3 (Storage Key-Time Monotonicity).
Given a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $K_{1}\subseteq K_{2}$:
@@
Intuitively, if $\textit{tx}_{1}\lesssim\textit{tx}_{2}$, by which we mean $t(\textit{tx}_{1})\leq t(\textit{tx}_{2})$ and $K(\textit{tx}_{1})\subseteq K(\textit{tx}_{2})$, then $\textit{tx}_{1}$ should cost no less than $\textit{tx}_{2}$. However, an attentive reader will observe that the former two are collectively equivalent to the latter (the proof and all subsequent omitted proofs can be found in the appendix):
@@
Property˜3 holds if and only if Properties˜1 and 2 hold.
Let us now take a moment to briefly evaluate why Properties˜1, 2 and 3 are not merely intuitive, but their violation can lead to harmful consequences and misaligned incentives: suppose $\textit{tx}_{1}\lesssim\textit{tx}_{2}$ and $T$ is a set of transactions containing neither of the two. If $\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}(\textit{tx}_{1})>\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}(\textit{tx}_{2})$, a user intending to submit $\textit{tx}_{1}$ might instead pad $\textit{tx}_{1}$ with unnecessary instructions and declare a larger access list to decrease the gas consumption. This would paradoxically reduce the gas usage and, assuming a reasonable TFM is used to compute transaction fees, also lower the transaction fee.
For any of the three properties above, we say that the property is strictly satisfied if for $\textit{tx}_{1}\not\simeq\textit{tx}_{2}$ the conclusion inequality holds strictly. Note that properties holding strictly are even more desirable with respect to the reasoning above: replacing a transaction with a “larger” one is then not only no better but actively worse. Unsurprisingly, Lemma˜3.1 also holds for the strict versions of the properties:
@@
Property˜3 holds strictly if and only if Properties˜1 and 2 hold strictly.
Next, we introduce another desirable monotonicity property, this time with a different emphasis: if two sets of transactions satisfy $T_{1}\subseteq T_{2}$, the transactions in $T_{1}$ should collectively consume no more gas than those in $T_{2}$ (Properties˜4 and 5). To formalize this, given a set of transactions $T$ and a subset $T^{\prime}\subseteq T$, write $\textsf{gas}_{T}(T^{\prime}):=\sum_{\textit{tx}^{\prime}\in T^{\prime}}\textsf{gas}_{T}(\textit{tx}^{\prime})$ for the total gas consumed by the transactions in $T^{\prime}$ when included in the set $T$ that constitutes a block.
@@
Similarly to before, we say that Property˜4 holds strictly if for $T_{1}\neq T_{2}$ the inequality in the conclusion holds strictly, which is desirable for reasons similar to those discussed above.
We now move on to more complex properties. Given two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2}),$ a transaction $\textit{tx}_{3}$ is their sequential composition (or, more simply, concatenation), if it executes the steps of $\textit{tx}_{1}$ followed by the steps of $\textit{tx}_{2}$. Note that, in this case $\textit{tx}_{3}\simeq(t_{1}+t_{2},K_{1}\cup K_{2})$. Our next property states that the concatenation of two transactions should consume no less gas than submitting them individually (Properties˜5 and 5). The two individual transactions perform the same actions as their concatenation, but not atomically, with no guarantee of their relative ordering or control over what happens between them. Naturally, enforcing atomicity and ordering limits the set of admissible concurrent schedules and requires storage keys to remain locked over longer continuous timespans. In particular, $\textit{tx}_{3}$ requires the storage keys in $K_{1}\cup K_{2}$ to be locked over a continuous span of $t_{1}+t_{2}$ units, while $\textit{tx}_{1}$ and $\textit{tx}_{2}$ submitted individually only require exclusive access to storage keys in $K_{1}$ for $t_{1}$ units and to storage keys in $K_{2}$ for $t_{2}$ units. Hence, the “bigger” transaction is at least as hard to schedule as its two constituent “parts” and should hence consume no less gas.
@@
We say that Property˜5 holds strictly if the inequality in the conclusion holds strictly, which is again more desirable than the basic version of the property.